OpenStudy (anonymous):

what the equation of a line in slope intercept form that runs through the points (4,5) and (-2,7)

OpenStudy (goldeverything):

do you know how to start out the problem

OpenStudy (anonymous):

sort of I know you do m= but I don't know how to do the rest

OpenStudy (goldeverything):

well try to take both of the coordinates and use them as a math problem

OpenStudy (goldeverything):

try subtracting them and tell me what you get

OpenStudy (anonymous):

how

OpenStudy (goldeverything):

\[(4-5)=\]

OpenStudy (goldeverything):

then do

OpenStudy (goldeverything):

sorry i messed up

OpenStudy (goldeverything):

\[(4-2)=\]

OpenStudy (goldeverything):

then you do

OpenStudy (goldeverything):

\[(5-7)=\]

OpenStudy (goldeverything):

now tell me what you get and take the first number as the x value and the second number as the y value

OpenStudy (goldeverything):

@kaylapietro have you gotten the awnser yet